MIT Services Status

All Systems Operational
General   Operational
Email   Operational
Network   Operational
Telephones   Operational
Academic Services   Operational
Administrative Services   Operational
Library Services   Operational
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Scheduled Maintenance
Due to scheduled Network Maintenance the MIT network may experience a brief interruption Tuesday 5/29/18 between 5:00am and 6:30am.

If you experience an outage lasting longer than posted, please contact the IS&T Service Desk at 617-253-1101 or servicedesk@mit.edu to open an incident report.
Posted on May 24, 15:02 EDT
As part of the MITnet Next Generation Project, the wired network IP address ranges for the listed buildings will be changed and placed behind the campus firewalls. Sporadic loss of network connectivity to/from the affected buildings during the maintenance window is expected.
Posted on May 17, 11:55 EDT
The Crashplan server will be unavailable Tuesday May 29th from 7p - 7:30p for routine maintenance. During this time, users will be unable to log in to the web console or initiate web restores. Restores initiated from desktop clients may experience brief interruptions during this time.
Posted on May 25, 11:56 EDT
As part of the MITnet Next Generation Project, the wired network IP address ranges for the listed buildings will be changed and placed behind the campus firewalls. Sporadic loss of network connectivity to/from the affected buildings during the maintenance window is expected.
Posted on May 17, 11:56 EDT
As part of the MITnet Next Generation Project, the wireless network IP address range will be changed. Sporadic loss of wireless (SSID = MIT SECURE) network connectivity during the maintenance window is expected.

The wireless (SSID = MIT) network will still be available during this maintenance window.

Current IP Range = 18.189.0.0/16
New IP Range = 18.40.0.0/16
Posted on May 24, 17:47 EDT
Past Incidents
May 26, 2018

No incidents reported today.

May 25, 2018
Completed - The scheduled maintenance has been completed.
May 25, 01:00 EDT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
May 24, 10:00 EDT
Scheduled - As part of the MITnet Next Generation Project, the wired network IP address ranges for the listed buildings will be changed and placed behind the campus firewalls. Sporadic loss of network connectivity to/from the affected buildings during the maintenance window is expected.
May 17, 11:53 EDT
May 23, 2018
Completed - The scheduled maintenance has been completed.
May 23, 14:00 EDT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
May 23, 10:00 EDT
Scheduled - As part of the MITnet Next Generation Project, the wired network IP address ranges for the listed buildings will be changed and placed behind the campus firewalls. Sporadic loss of network connectivity to/from the affected buildings during the maintenance window is expected.
May 17, 11:51 EDT
May 22, 2018
Completed - The scheduled maintenance has been completed.
May 22, 16:00 EDT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
May 22, 10:00 EDT
Scheduled - As part of the MITnet Next Generation Project, the wired network IP address ranges for the listed buildings will be changed and placed behind the campus firewalls. Sporadic loss of network connectivity to/from the affected buildings during the maintenance window is expected.
May 17, 11:50 EDT
Resolved - This incident has been resolved.
May 22, 12:04 EDT
Update - While the OC11 firewalls have recovered we are aware of several services that still have not recovered. We are working on stablizing those services as we learn about them.
May 18, 14:36 EDT
Monitoring - IS&T performed an emergency software upgrade to the OC11 data center firewalls. All services were restored as of 11:30am. IS&T will continue to monitor the situation.
May 18, 11:51 EDT
Investigating - IS&T is aware of another incident affecting core services and is investigating the issue.
May 18, 09:56 EDT
May 21, 2018
Completed - The scheduled maintenance has been completed.
May 21, 22:00 EDT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
May 21, 10:00 EDT
Scheduled - As part of the MITnet Next Generation Project, the wired network IP address ranges for the listed buildings will be changed and placed behind the campus firewalls. Sporadic loss of network connectivity to/from the affected buildings during the maintenance window is expected.
May 17, 11:48 EDT
May 20, 2018

No incidents reported.

May 19, 2018

No incidents reported.

May 18, 2018
Resolved - Core services were interrupted due to a firewall hardware crash at the OC11 data center at approximately 8:40am. IS&T network engineers failed over to redundant hardware at approximately 9:10am. All core services have been restored. We apologize for the inconvenience.
May 18, 09:21 EDT
Update - IS&T has resolved the issue.
May 18, 09:20 EDT
Investigating - IS&T has noticed that many services are currently intrrupted we are currently investigating the scope and impact.
May 18, 09:00 EDT
May 17, 2018
Completed - The scheduled maintenance has been completed.
May 17, 22:00 EDT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
May 17, 18:01 EDT
Scheduled - As part of the MITnet Next Generation Project, the wired network IP address ranges for the listed buildings will be changed and placed behind the campus firewalls. Sporadic loss of network connectivity to/from the affected buildings during the maintenance window is expected.
May 17, 11:47 EDT
Completed - The scheduled maintenance has been completed.
May 17, 22:00 EDT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
May 17, 18:01 EDT
Scheduled - As part of the MITnet Next Generation Project, the wired network IP address ranges for the listed buildings will be changed. Sporadic loss of network connectivity to/from the affected buildings during the maintenance window is expected.
May 15, 11:24 EDT
Resolved - This incident has been resolved.
May 17, 14:10 EDT
Investigating - There was an issue with the feed last night from Roles into the Warehouse which is causing users to have limited access to warehouse objects. The warehouse is currently looking into the problem and will have it resolved as soon as possible.
May 17, 11:00 EDT
May 16, 2018

No incidents reported.

May 15, 2018
Completed - The scheduled maintenance has been completed.
May 15, 20:00 EDT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
May 15, 08:00 EDT
Scheduled - Maintenance to the MIT Regional Optical Network requires modifying optical equipment at Albany and OC40 (MGHPCC). Due to optical ring redundancy, and the manner in which the maintenance will be conducted, no outage or loss of connectivity is expected.
May 3, 16:02 EDT
Completed - The scheduled maintenance has been completed.
May 15, 19:00 EDT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
May 15, 18:00 EDT
Update - We will be undergoing scheduled maintenance during this time.
May 14, 12:18 EDT
Scheduled - IS&T will be upgrading the VoIP network in Building E19. Phones will need to reboot once the change is made.
May 11, 16:08 EDT
Completed - The scheduled maintenance has been completed.
May 15, 18:00 EDT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
May 15, 10:01 EDT
Scheduled - As part of the MITnet Next Generation Project, the wired network IP address ranges for the listed buildings will be changed and placed behind the campus firewalls. Sporadic loss of network connectivity to/from the affected buildings during the maintenance window is expected.
May 1, 12:26 EDT
Resolved - This incident has been resolved.
May 15, 13:38 EDT
Identified - IS&T is aware of a certificate expiration issue with reports.mit.edu, impacting Cognos reporting. We are working on a resolution and will post updates here. We apologize for the inconvenience.
May 15, 13:03 EDT
Completed - The scheduled maintenance has been completed.
May 15, 08:00 EDT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
May 15, 05:00 EDT
Scheduled - Due to a scheduled power outage the MIT wired / BMS networks and ISDN telephones will be down on the first floor and isolated pockets of the basement for the duration of the outage starting Tuesday May 15th at 5am until 8am.

If you experience an outage lasting longer than posted, please contact the IS&T Service Desk at 617-253-1101 or servicedesk@mit.edu to open an incident report.
May 10, 16:14 EDT
May 14, 2018
Resolved - This incident has been resolved.
May 14, 11:44 EDT
Update - Connectivity to the remaining houses has been restored.

IS&T is now verifying any related issues or work has been completed.
May 13, 20:21 EDT
Update - Connectivity to the FSILG houses off the 120 Baystate Rd. fiber run (BTP, TX, PDT, SAE, SH, AEPi, LCA) has been restored. IS&T is still confirming with the fiber provider resolution is complete.

The FSILG houses off the 479 Commonwealth Ave fiber run (AP, FH, PSK, SN, ACO, SK) are still affected. More in depth tests of the new fiber circuit for this run is in progress.
May 13, 19:08 EDT
Update - The fiber provider has now made further repairs beyond the initial required splicing. While some repairs have been completed, multiple customers (including MIT/FSILG) are still affected. The fiber provider is currently conducting local on-site (FSILG) tests to resolve the issue.
May 13, 15:28 EDT
Update - We continue to work with our fiber optic contractor to restore service to the off-campus houses affected by the fiber cut in Boston. The contractor is dispatching technicians to the next hop after the cut in order to do further diagnosis.
May 13, 10:34 EDT
Update - Fiber splicing has now been completed, but alarms are still being cleared by the fiber provider. They are currently unable to provide an estimate for the full time of repair.
May 12, 21:25 EDT
Update - Our fiber contractor has indicated that splicing is well under way. They estimate restoration of services at approximately 2pm today (Saturday 10/12).
May 12, 11:03 EDT
Update - We've been informed that our fiber contractor was granted access to closely assess the damaged cable. One side of the break has been prepared for spliicing, and work will begin on the other side shortly. Splicing is expected to take place at about noon today, Saturday 5/12.
May 12, 07:32 EDT
Update - An update received from our fiber provider at 3pm has identified this as a major fiber optic cable cut. They continue to work on repair. MIT is one among many affected customers.
May 11, 16:29 EDT
Identified - The following FSILG locations are experiencing a network outage: BTP, TX, PDT, LCA, SH, BTP-119, AEPI, SAE, AP, PSK, SK, ACO, SN, FH, and ZBT.

Our fiber provider has identified a fiber cut and is investigating.
May 11, 10:18 EDT
May 13, 2018
Completed - The scheduled maintenance has been completed.
May 13, 23:01 EDT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
May 11, 17:00 EDT
Scheduled - Request Tracker will be undergoing infrastructure maintenance. Short ( > 15 mins ) latencies might be experienced a couple of time during this maintenance. If you have any questions or concerns, please call the Service Desk at 617-253-1101 or email us at servicedesk@mit.edu.
May 11, 14:49 EDT
May 12, 2018
Completed - The scheduled maintenance has been completed.
May 12, 19:10 EDT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
May 12, 11:10 EDT
Scheduled - We received late notice of scheduled electric power work at Building EE19 (1 Rogers St).
May 12, 11:07 EDT
Completed - The scheduled maintenance has been completed.
May 12, 01:00 EDT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
May 12, 00:00 EDT
Scheduled - Quick Base will be switching the active instance of the Quick Base platform from their data center in Denver, CO to that in Las Vegas, NV.

To provide a highly available platform, Quick Base is hosted in two geographically diverse data centers and can run actively in only one location at a time with the other running in standby mode.

The vendor is planning to switch between their data centers on Saturday, May 12th, between 12:00 AM and 1:00 AM Eastern US Time. During this switchover, Quick Base will be down for no more than 1 hour. Users will not need to do anything before May 12th. Additional details are available on Quick Base Community: http://stats.quickbase.com/a0100v0iyB0W008NEAt9qUi.

As part of this switchover, there is no expected change in how the platform functions or performs and Quick Base will uphold the platform's high availability, disaster recovery, and security and compliance standards. For more on Quick Base security and compliance standards, see: http://stats.quickbase.com/Y0iq909tU00yW001vAB0FNi
May 1, 10:40 EDT